# Rémi A. Chou and Matthieu R. Bloch

### v1

@@ Abstract
We consider multi-user commitment models that capture the problem of enabling multiple bidders to simultaneously submit auctions to verifiers while ensuring that i) verifiers do not obtain information on the auctions until bidders reveal them at a later stage; and, ii) bidders cannot change their auction once committed. Specifically, we assume that bidders and verifiers have access to a noiseless channel as well as a noisy multiple-access channel or broadcast channel, where inputs are controlled by the bidders and outputs are observed by verifiers. In the case of multiple bidders and a single verifier connected by a non-redundant multiple-access channel, we characterize the commitment capacity region when bidders are not colluding. When the bidders are colluding, we derive an achievable region and a tight converse for the sum rate. In both cases our proposed achievable commitment schemes are constructive. In the case of a single bidder and multiple verifiers connected by a non-redundant broadcast channel, in which verifiers could drop out of the network after auctions are committed, we also characterize the commitment capacity region. Our results demonstrate how commitment schemes can benefit from multi-user protocols, and develop resilience when some verifiers may become unavailable.
@@ I Introduction
Commitment without the need for a trusted third party can be traced back to Blum’s coin-flipping problem [2]. More generally, a two-party commitment problem involves a bidder, Alice, and a verifier, Bob, and operates in two phases. In the first phase, called the commit phase, Alice sends Bob information to commit to a message $M$, representing a bid in an auction that must remain concealed from Bob. In the second phase, called the reveal phase, Alice reveals a message $M^{\prime}$ to Bob, who must determine whether $M^{\prime}$ is the message that Alice committed to in the commit phase. The protocol must be binding in the sense that, in the reveal phase, Alice cannot make Bob believe that she committed to a message $M^{\prime}\neq M$. It is well-known that information-theoretic concealment guarantees cannot be achieved over noiseless communication channels, e.g., [3]. However, when a noisy channel is available as a resource, both concealment and binding requirements can be obtained under information-theoretic guarantees, i.e., when Alice and Bob are not assumed to be computationally limited, for some class of noisy channels called non-redundant [4].
@@
Most of the literature on information-theoretic commitments focuses on two-party scenarios that involve a single bidder and a single verifier, e.g., [5, 3, 4, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15], characterizing commitment capacity under increasingly complex channel models. We study here instead two specific multi-user commitment settings: one in which a verifier interacts with $L$ bidders, each committing to individual messages, and another in which a bidder commits to $B$ verifiers, some of whom may drop out after the commit phase. The motivation for our first setting is to explore whether a multi-bidder protocol can outperform single-bidder protocols and time-sharing when multiple bidders wish to commit to individual messages. Our motivation to consider multiple verifiers in our second setting, is to ensure positive commitment rates even if a verifier drops out of the network after the commit phase, which would be impossible with a single verifier.
@@
In our first setting, $L$ bidders and the verifier have access to a noiseless public communication channel and a noisy discrete memoryless multiple-access channel with $L$ inputs. Each input of the multiple-access channel is controlled by a distinct bidder and the verifier observes the output of the channel. Similar to a two-party setting, the protocol consists of a commit phase and a reveal phase. Here, the concealment requirement is that the verifier must not learn, in an information-theoretic sense, information about any message of any bidder after the commit phase. The protocol must also be information-theoretically binding in the sense that, during the reveal phase, a bidder cannot make the verifier believe that it committed to another message than the one committed to in the commit phase. For this setting, we consider both the cases of colluding and non-colluding bidders. The non-colluding bidders case corresponds to a scenario in which the bidders do not trust each other and do not want to exchange information with one another. For instance, this would be the case when the bidders commit to messages sent to an auctioneer. Under a non-redundancy condition on the multiple-access channel, we derive the capacity region for the non-colluding bidders case, and an achievable region and the sum-rate capacity for the colluding bidders case. In both cases, our achievability scheme is constructive and relies on distributing hashing with two-universal hash functions [16] for the concealment guarantees. The bindingness of our achievability scheme hinges on the non-redundancy property of the multiple access channel akin to the two-party commitment in [4]. The characterization of the sum-rate capacity relies on the polymatroidal properties of our achievability region. Additionally, we demonstrate through a numerical example that, for some channels, using a multi-bidder protocol can outperform using single-bidder protocols, e.g., [4], and time-sharing.
@@
The remainder of the paper is organized as follows. After a review of notation in Section II, we develop an auxiliary result in Section III, offering a sufficient condition to guarantee non-redundancy and an alternative characterization of non-redundant channels with input alphabet of cardinality at most three. We formally introduce our multiuser commitment models in Section IV along with the associated results. We delegate the proofs to Section V and Section VI to streamline the presentation. Finally, we provide concluding remark in Section VII.
@@ II Notation
[…]
@@ III Non-Redundant Channels
[4] introduced the concept of non-redundant channel as follows. A DMC $(\mathcal{X},\mathcal{Y},W)$ is non-redundant if
$\exists\eta>0,\forall x\in\mathcal{X},\forall p\in\mathcal{P}(\mathcal{X})\text{ s.t. }p(x)=0,\|W_{x}-W\circ p\|\geqslant\eta.$ (1)
@@ A DMC $(\mathcal{X},\mathcal{Y},W)$ is non-redundant if
|$\absolutevalue{f(p_{1})-f(p_{2})}$ | =|$=\absolutevalue{\|W_{x}-W\circ p_{1}\|-\|W_{x}-W\circ p_{2}\|}$ |
@@
so that $f_{x}$ is continuous on $\Delta(\mathcal{X}\setminus\{x\})$. Since $\Delta(\mathcal{X}\setminus\{x\})$ is a subset of |$\mathbb{R}^{\absolutevalue{\mathcal{X}}-1}$ that is closed and bounded, it is compact and there exists $p_{x}\in\Delta(\mathcal{X}\setminus\{x\})$ such that
@@
If |$\absolutevalue{\mathcal{X}}\leqslant 3$, a DMC $(\mathcal{X},\mathcal{Y},W)$ is non-redundant if and only if
@@
which have disjoint support, and note that $\sum_{u\in\mathcal{S}}(p(u)-q(u))=\sum_{u\in\mathcal{S}^{c}}(q(u)-p(u))$. Then, we have
@@
$\Leftrightarrow\forall y\in\mathcal{Y},\sum_{x\in\mathcal{S}}W(y|x)(p(x)-q(x))=\sum_{x\in\mathcal{S}^{c}}W(y|x)(q(x)-p(x))$
$\Leftrightarrow\forall y\in\mathcal{Y},\sum_{x\in\mathcal{S}}W(y|x)\tilde{p}(x)=\sum_{x\in\mathcal{S}^{c}}W(y|x)\tilde{q}(x).$
Since |$\absolutevalue{\mathcal{X}}\leqslant 3$, either $|\mathcal{S}|=1$ or $|\mathcal{S}^{c}|=1$ and we can assume without loss of generality that $\mathcal{S}=\{x^{*}\}$ and $\tilde{p}(x^{*})=1$. Hence, $W_{x^{*}}=\sum_{x\in\mathcal{X}\setminus\{x^{*}\}}W(y|x)\tilde{q}(x)$ and the channel is redundant as per (2). ∎
Proposition 2 provides an alternative characterization of the non-redundant condition but the result cannot be extended to the case $|\mathcal{X}|>3$. Indeed, the channel with |$\absolutevalue{\mathcal{X}}=4$, |$\absolutevalue{\mathcal{Y}}=3$, and
@@ IV-A Multi-Bidder Single-Verifier Model and Results
$\forall x_{\mathcal{L}}\in\mathcal{X}_{\mathcal{L}},\forall p_{X_{\mathcal{L}}}\in\mathcal{P}(\mathcal{X}_{\mathcal{L}})\text{ s.t. }p_{X_{\mathcal{L}}}(x_{\mathcal{L}})=0,W_{x_{\mathcal{L}}}\neq\sum_{x^{\prime}_{\mathcal{L}}\in\mathcal{X}_{\mathcal{L}}}p_{X_{\mathcal{L}}}(x^{\prime}_{\mathcal{L}})W_{x^{\prime}_{\mathcal{L}}}.$ (8)
@@ IV-A1 Colluding bidders
    Reveal phase: The bidders reveal $(a_{\mathcal{L}},S)$ to the verifier. The verifier performs a test $\beta(V,a_{\mathcal{L}},S)$ that returns $1$ if the sequence $a_{\mathcal{L}}$ is accepted and 00 otherwise.
@@
A rate-tuple $(R_{\ell})_{\ell\in\mathcal{L}}$ is achievable if there exists a sequence of $((2^{n{R_{\ell}}})_{{\ell}\in\mathcal{L}},n)$ commitment schemes such that for any $\tilde{S}\in\mathcal{R}$, $a_{\mathcal{L}},a_{\mathcal{L}}^{\prime}\in\mathcal{A}_{\mathcal{L}}$ such that $a_{\mathcal{L}}\neq a^{\prime}_{\mathcal{L}}$,
[…]
@@
(9) means that the verifier accepts the revealed sequences $a_{\mathcal{L}}$ as long as the bidders are honest. (10) ensures that the verifier gains no information about the committed sequences $a_{\mathcal{L}}$ prior to the reveal phase. (11) ensures that the bidders cannot convince the verifier to accept sequences other than those initially committed to.
Our main results is a partial characterization of the capacity region for colluding bidders. In the following, for any $\mathcal{T}\subseteq\mathcal{L}$, we write the sum-rate of the bidders in $\mathcal{T}$ as $R_{\mathcal{T}}\triangleq\sum_{{\ell}\in\mathcal{T}}R_{\ell}$.
@@
The following example demonstrates that, for some channels, having a multi-bidder coding scheme, rather than doing time-sharing with single-bidder coding schemes as in [4], can increase the sum-rate of the bidders.
@@
The transition probability matrix $W_{i}$, $i\in\{1,2\}$, corresponds to the situation in which only Bidder $(1-i)$ uses the channel $W$ and $X_{i}$ is constant and equal to 00, i.e., when the distribution of $X_{i}$ is $p_{X_{i}}^{\star}$ defined by $(p_{X_{i}}^{\star}(0),p_{X_{i}}^{\star}(1))\triangleq(1,0)$. For the problem of commitment as described in Definition 1, let $\Sigma(W)$ be the sum-rate capacity for the channel $W$ and $C(W_{i})$, $i\in\{1,2\}$, be the commitment capacity for the channel $W_{i}$. For convenience, we write $H_{p}(X_{1}X_{2}|Y)$ to indicate that $(X_{1},X_{2})$ is distributed according to $p$. Then, from Theorem 1, we have
@@
$\Sigma(W)>1.9647\textnormal{ bits/channel use}>0.9512\textnormal{ bits/channeluse}>\max(C(W_{1}),C(W_{2})).$
@@ IV-A2 Non-colluding bidders
    Reveal phase: Bidder $\ell\in\mathcal{L}$ reveals $(a_{\ell},S_{\ell})$ to the verifier. For each ${\ell}\in\mathcal{L}$, the verifier performs a test $\beta_{\ell}(V_{\ell},a_{\ell},S_{\ell})$, ${\ell}\in\mathcal{L}$, that returns $1$ if the sequence $a_{\ell}$ is accepted and 00 otherwise.
@@
A rate-tuple $(R_{\ell})_{\ell\in\mathcal{L}}$ is achievable if there exists a sequence of $((2^{n{R_{\ell}}})_{{\ell}\in\mathcal{L}},n)$ commitment schemes such that for any ${\ell}\in\mathcal{L}$, $\tilde{S}_{\ell}\in\mathcal{R}_{\ell}$, $a_{\mathcal{L}},a_{\mathcal{L}}^{\prime}\in\mathcal{A}_{\mathcal{L}}$ such that $a_{\mathcal{L}}\neq a^{\prime}_{\mathcal{L}}$,
[…]
@@
Similar to the scenario with colluding bidders, the following example illustrates that, for some channels, a multi-bidder coding scheme can again outperform a time-sharing approach that utilizes single-bidder coding schemes as in [4].
@@
$\Sigma^{\textup{I}}(W)>1.9645\textnormal{ bits/channel use}>0.9512\textnormal{bits/channel use}>\max(C(W_{1}),C(W_{2})).$

### v2

@@ Abstract
We consider multi-user commitment models that capture the problem of enabling multiple bidders to simultaneously submit auctions to verifiers while ensuring that i) verifiers do not obtain information on the auctions until bidders reveal them at a later stage; and, ii) bidders cannot change their auction once committed. Specifically, we assume that bidders and verifiers have access to a noiseless channel as well as a noisy multiple-access channel or broadcast channel, where inputs are controlled by the bidders and outputs are observed by verifiers. In the case of multiple bidders and a single verifier connected by a non-redundant multiple-access channel, we characterize the commitment capacity region when bidders are not colluding. When the bidders are colluding, we derive an achievable region and a tight converse for the sum rate. In both cases our proposed achievable commitment schemes are constructive. In the case of a single bidder and multiple verifiers connected by a non-redundant broadcast channel, in which verifiers could drop out of the network after auctions are committed, we also characterize the commitment capacity. Our results demonstrate how commitment schemes can benefit from multi-user protocols, and develop resilience when some verifiers may become unavailable.
@@ I Introduction
Commitment without the need for a trusted third party can be traced back to Blum’s coin-flipping problem [2]. More generally, a two-party commitment problem involves a bidder, Alice, and a verifier, Bob, and operates in two phases. In the first phase, called the commit phase, Alice sends Bob information to commit to a message $M$, representing a bid in an auction that must remain concealed from Bob. In the second phase, called the reveal phase, Alice reveals a message $M^{\prime}$ to Bob, who must determine whether $M^{\prime}$ is the message that Alice committed to in the commit phase. The protocol must be binding in the sense that, in the reveal phase, Alice cannot make Bob believe that she committed to a message $M^{\prime}\neq M$. It is well-known that information-theoretic concealment guarantees cannot be achieved over noiseless communication channels and in the absence of additional resources, e.g., [3]. However, when a noisy channel is available as a resource, both concealment and binding requirements can be obtained under information-theoretic guarantees, i.e., when Alice and Bob are not assumed to be computationally limited, for some class of noisy channels called non-redundant [4].
@@
Most of the literature on information-theoretic commitments focuses on two-party scenarios that involve a single bidder and a single verifier, e.g., [5, 3, 4, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15], characterizing commitment capacity under increasingly complex channel models. We study here instead two specific multi-user commitment settings: one in which a verifier interacts with $L$ bidders, each committing to individual messages, and another in which a bidder commits to $B$ verifiers, some of whom may drop out after the commit phase. The motivation for our first setting is to explore whether a multi-bidder protocol can outperform single-bidder protocols and time-sharing when multiple bidders wish to commit to individual messages. Our motivation to consider multiple verifiers in our second setting, is to ensure positive commitment rates even if a verifier drops out of the network after the commit phase, which would be impossible with a single verifier. We note that extensions of the single-user commitment problems can also be found in the computer science literature. For instance, vector commitment [16] addresses the problem of committing ordered sequences that can later be opened in specific positions. The key contribution therein is to create concise commitments, for which the commitment size is independent of the committed vector length. While the conciseness of the proposed vector commitment bears conceptual similarities with the notion of commitment capacity, the security is computational and does not rely on a noisy channel. Perhaps more closely related to the present work, the notion of commitment with multiple committers and verifiers under information-theoretic security has also been explored in the absence of a noisy channel in [17], when each pair of parties has access to a secure private channel and the majority of parties are honest. Because [17] does not exploit noisy channels, the coding mechanisms enabling commitment are fundamentally different.
@@
In our first setting, $L$ bidders and the verifier have access to a noiseless public communication channel and a noisy discrete memoryless multiple-access channel with $L$ inputs. Each input of the multiple-access channel is controlled by a distinct bidder and the verifier observes the output of the channel. Similar to a two-party setting, the protocol consists of a commit phase and a reveal phase. Here, the concealment requirement is that the verifier must not learn, in an information-theoretic sense, information about any message of any bidder after the commit phase. The protocol must also be information-theoretically binding in the sense that, during the reveal phase, a bidder cannot make the verifier believe that it committed to another message than the one committed to in the commit phase. For this setting, we consider both the cases of colluding and non-colluding bidders. The non-colluding bidders case corresponds to a scenario in which the bidders do not trust each other and do not want to exchange information with one another. For instance, this would be the case when the bidders commit to messages sent to an auctioneer. Note that the case of colluding bidders may not be suitable for an auction, as bidders are allowed to share their committed sequences. The motivation behind our colluding bidders setting is to enable bulk commitment, for example, in scenarios where each bidder represents a data source, and all data sources belong to the same entity. In such cases, data sources commit their values to the verifier collectively. Example 1 illustrates that, for some channels, jointly committing across data sources can be more advantageous than separately committing. Under a non-redundancy condition on the multiple-access channel, we derive the capacity region for the non-colluding bidders case, and an achievable region and the sum-rate capacity for the colluding bidders case. In both cases, our achievability scheme is constructive and relies on distributing hashing with two-universal hash functions [18] for the concealment guarantees. The bindingness of our achievability scheme hinges on the non-redundancy property of the multiple access channel akin to the two-party commitment in [4]. The characterization of the sum-rate capacity relies on the polymatroidal properties of our achievability region. Additionally, we demonstrate through a numerical example that, for some channels, using a multi-bidder protocol can outperform using single-bidder protocols, e.g., [4], and time-sharing.
@@
We emphasize that the analysis of multi-user commitment goes beyond the single-user case on several accounts: i) the converse proof requires careful considerations to properly handle Markov chains; ii) the key generation process involves a multi-user version of the hash lemma; and, iii) we invoke sub-modularity arguments in some of our achievability proofs.
@@
The remainder of the paper is organized as follows. After a review of notation in Section II, we develop an auxiliary result in Section III, offering a sufficient condition to guarantee non-redundancy and an alternative characterization of non-redundant channels with input alphabet of cardinality at most three. We formally introduce our multiuser commitment models in Section IV along with the associated results. We delegate the proofs to Section V and Section VI. Finally, we provide concluding remark in Section VII.
@@ II Notation
For $a,b\in\mathbb{R}$, define $\llbracket a,b\rrbracket\triangleq[\lfloor a\rfloor,\lceil b\rceil]\cap\mathbb{N}$. Unless specified otherwise, random variables and their realizations are denoted by uppercase and corresponding lowercase letters, respectively, e.g., $x$ is a realization of the random variable $X$. A Discrete Memoryless Channel (DMC) with input alphabet $\mathcal{X}$, output alphabet $\mathcal{Y}$, and transition probability $W$ is denoted by $(\mathcal{X},\mathcal{Y},W)$. Additionally, for $x\in\mathcal{X}$, $W_{x}$ denotes the output distribution of the channel when the input is $x$. The probability simplex for distributions defined over the set $\mathcal{X}$ is denoted by $\mathcal{P}(\mathcal{X})$. For a DMC $(\mathcal{X},\mathcal{Y},W)$ and a distribution $p\in\mathcal{P}(\mathcal{X})$, $W\circ p$ denotes the distribution of the channel output when the input is distributed according to $p$. $\|\cdot\|$ denotes the $\ell_{1}$-norm. For two distributions $p,q\in\mathcal{P}(\mathcal{X})$, the variational distance between $p$ and $q$ is denoted by $\mathbb{V}(p,q)\triangleq\frac{1}{2}\|p-q\|$. The indicator function is denoted by $\mathds{1}\{\omega\}$, which is equal to $1$ if the predicate $\omega$ is true and 0 otherwise.
@@ III Non-Redundant Channels
The concept of non-redundant channel was introduced in [4]. A DMC $(\mathcal{X},\mathcal{Y},W)$ is non-redundant if
$\exists\eta>0,\forall x\in\mathcal{X},\forall p\in\mathcal{P}(\mathcal{X})\text{ s.t. }p(x)=0,$
$\|W_{x}-W\circ p\|\geqslant\eta.$ (1)
@@ A DMC $(\mathcal{X},\mathcal{Y},W)$ is non-redundant if
$|\displaystyle\absolutevalue{f(p_{1})-f(p_{2})}$ $|\displaystyle=\absolutevalue{\|W_{x}-W\circ p_{1}\|-\|W_{x}-W\circ p_{2}\|}$
@@
[…]
@@
[…]
@@
which have disjoint support, and note that
$\sum_{u\in\mathcal{S}}(p(u)-q(u))=\sum_{u\in\mathcal{S}^{c}}(q(u)-p(u)).$
Then, we have
@@
$\Leftrightarrow\forall y\in\mathcal{Y},\textstyle\sum_{x\in\mathcal{S}}W(y|x)(p(x)-q(x))$
$\phantom{-------}=\textstyle\sum_{x\in\mathcal{S}^{c}}W(y|x)(q(x)-p(x))$
$\Leftrightarrow\forall y\in\mathcal{Y},\textstyle\sum_{x\in\mathcal{S}}W(y|x)\tilde{p}(x)=\textstyle\sum_{x\in\mathcal{S}^{c}}W(y|x)\tilde{q}(x).$
[…]
@@ IV-A Multi-Bidder Single-Verifier Model and Results
$\forall x_{\mathcal{L}}\in\mathcal{X}_{\mathcal{L}},\forall p_{X_{\mathcal{L}}}\in\mathcal{P}(\mathcal{X}_{\mathcal{L}})\text{ s.t. }p_{X_{\mathcal{L}}}(x_{\mathcal{L}})=0,$
$W_{x_{\mathcal{L}}}\neq\sum_{x^{\prime}_{\mathcal{L}}\in\mathcal{X}_{\mathcal{L}}}p_{X_{\mathcal{L}}}(x^{\prime}_{\mathcal{L}})W_{x^{\prime}_{\mathcal{L}}}.$ (8)
@@ IV-A1 Colluding bidders
    Reveal phase: The bidders reveal $(a_{\mathcal{L}},S)$ to the verifier. The verifier performs a test $\beta(V,a_{\mathcal{L}},S)$ that returns $1$ if the sequence $a_{\mathcal{L}}$ is accepted and 0 otherwise.
@@
A rate-tuple $(R_{\ell})_{\ell\in\mathcal{L}}$ is achievable if there exists a sequence of $((2^{n{R_{\ell}}})_{{\ell}\in\mathcal{L}},n)$ commitment schemes such that for any $\tilde{S}\in\mathcal{R}$, $a_{\mathcal{L}},\tilde{a}_{\mathcal{L}}\in\mathcal{A}_{\mathcal{L}}$ such that $a_{\mathcal{L}}\neq\tilde{a}_{\mathcal{L}}$,
$\lim_{n\to\infty}\mathbb{P}[\beta(V,a_{\mathcal{L}},S)=0]$ $=0,\text{ (correctness) }$ (9)
$\lim_{n\to\infty}I(A_{\mathcal{L}};V)$ $=0,\text{ (concealment)}$ (10)
@@
Equation (9) means that the verifier accepts the revealed sequences $a_{\mathcal{L}}$ as long as the bidders are honest. Equation (10) ensures that the verifier gains no information about the committed sequences $a_{\mathcal{L}}$ prior to the reveal phase. Equation (11) ensures that the bidders cannot convince the verifier to accept sequences other than those initially committed to.
Our main result is a partial characterization of the capacity region for colluding bidders. In the following, for any $\mathcal{T}\subseteq\mathcal{L}$, we write the sum-rate of the bidders in $\mathcal{T}$ as $R_{\mathcal{T}}\triangleq\sum_{{\ell}\in\mathcal{T}}R_{\ell}$.
@@
The following example demonstrates that, for some channels, having a multi-bidder coding scheme, rather than doing time-sharing with single-bidder coding schemes [4], can increase the sum-rate of the bidders.
@@
The transition probability matrix $W_{i}$, $i\in\{1,2\}$, corresponds to the situation in which only Bidder (3i)(3-i) uses the channel $W$ and $X_{i}$ is constant and equal to 0, i.e., when the distribution of $X_{i}$ is $p_{X_{i}}^{\star}$ defined by $(p_{X_{i}}^{\star}(0),p_{X_{i}}^{\star}(1))\triangleq(1,0)$. For the problem of commitment as described in Definition 1, let $\Sigma(W)$ be the sum-rate capacity for the channel $W$ and $C(W_{i})$, $i\in\{1,2\}$, be the commitment capacity for the channel $W_{i}$. For convenience, we write $H_{p}(X_{1}X_{2}|Y)$ to indicate that $(X_{1},X_{2})$ is distributed according to $p$. Then, from Theorem 1, we have
@@
$\Sigma(W)$ $>1.9647\textnormal{ bits/channel use}$
$>0.9512\textnormal{ bits/channel use}$
$>\max(C(W_{1}),C(W_{2})).$
@@ IV-A2 Non-colluding bidders
    Reveal phase: Bidder $\ell\in\mathcal{L}$ reveals $(a_{\ell},S_{\ell})$ to the verifier. For each $\ell$$\in\mathcal{L}$, the verifier performs a test $\beta_{\ell}(V_{\ell},a_{\ell},S_{\ell})$, $\ell$$\in\mathcal{L}$, that returns $1$ if the sequence $a_{\ell}$ is accepted and 0 otherwise.
@@
A rate-tuple $(R_{\ell})_{\ell\in\mathcal{L}}$ is achievable if there exists a sequence of $((2^{n{R_{\ell}}})_{{\ell}\in\mathcal{L}},n)$ commitment schemes such that for any $\ell$$\in\mathcal{L}$, $\tilde{S}_{\ell}\in\mathcal{R}_{\ell}$, $a_{\mathcal{L}},\tilde{a}_{\mathcal{L}}\in\mathcal{A}_{\mathcal{L}}$ such that $a_{\mathcal{L}}\neq\tilde{a}_{\mathcal{L}}$,
$\lim_{n\to\infty}\mathbb{P}[\beta_{\ell}(V_{\ell},a_{\ell},S_{\ell})=0]$ $=0,\text{ (correctness) }$
$\lim_{n\to\infty}I(A_{\mathcal{L}};V_{\mathcal{L}})$ $=0,\text{ (concealment)}$
@@
Similar to the scenario with colluding bidders, the following example illustrates that, for some channels, a multi-bidder coding scheme can again outperform a time-sharing approach that utilizes single-bidder coding schemes [4].
@@
$\Sigma^{\textup{I}}(W)$ $>1.9645\textnormal{ bits/channel use}$
$>0.9512\textnormal{ bits/channel use}$
$>\max(C(W_{1}),C(W_{2})).$